OpenStudy (anonymous):

What is the value of the missing exponent? (4x^4)^?=16x^8

OpenStudy (anonymous):

2

OpenStudy (anonymous):

(4x4)t = 16x8 16t = 16x8 t = (16x8)/16 t = 8

OpenStudy (anonymous):

4^2=16 (x^4)^2=x^8

jhonyy9 (jhonyy9):

2 is sure

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2